# Unbundle-Rewrite-Rebundle: Runtime Detection and Rewriting of Privacy-Harming Code in JavaScript Bundles

### v2

@@ 6. Related Work
This work contributes to and builds on an enormous body of research relevant to web privacy, tracking, and content blocking. In this section, we highlight existing research that relates to the our framework’s design.
@@
This research is closely tied to a broad domain of research investigating the advantages, effectiveness, and responses to filter list-based content blocking. Note that existing filter lists (e.g., EasyList (EasyList Authors, 2023), EasyPrivacy (EasyPrivacy Authors, 2023)) are developed manually and community-maintained. Merzdovnik et al. (Merzdovnik et al., 2017) performed a large-scale study that highlighted the effectiveness of filter list-based browser extensions and also indicated that these tools often lead to a decrease in overall CPU usage, even when considering their own overhead. Gervais et al. (Gervais et al., 2017) quantified the privacy gain from ad-blockers and discovered that these tools can reduce interactions with third-party entities by up to 40% with default settings.
However, other studies have pointed out significant inefficiencies in filter lists. Snyder et al. (Snyder et al., 2020) highlighted the abundance of “dead-weight” rules that offer no discernible benefits in popular lists. Similarly, Alrizah et al. (Alrizah et al., 2019) found that popular lists contain a large number of false positives which can take two or three months to be discovered.
@@
Numerous studies have developed approaches to either help automatically generate filter list rules or develop alternative methods to block privacy harming content. AdGraph (Iqbal et al., 2020) created a graph-based machine learning approach that used features like URL length and origin to differentiate between benign and privacy harming resources. Bhagavatula et al. (Bhagavatula et al., 2014) also used URL-based features to train a machine learning model for resource filtering. Chen et al. (Chen et al., 2021) used filter lists as ground truth and developed behavioral signatures based on the JavaScript event loop, while Sun et al. (Sun et al., 2021) classified JavaScript execution based on Web API calls. Le et al. (Le et al., 2023) developed a reinforcement learning framework that generates filter list rules specific to a site of interest and showed that their approach was comparable in visual breakage to manually-created filter lists.
@@
Prior work that perform static or dynamic analysis on JavaScript to create “pre-filters” (Fass et al., 2019b) for malicious scripts, detect malicious scripts that camouflage as benign scripts (Fass et al., 2019a), and detect scripts that evade detection by adopting various obfuscation strategies (Fass et al., 2018). In their analysis of the Top 10K sites in the Alexa list, Moog et al. (Moog et al., 2021) found that 90% sites contained a minified or obfuscated script.
More relevant to our work, Rack and Staicu (Rack and Staicu, 2023b) presented a method for detecting and partially reverse engineering bundles. However, unlike the automated approach we present in § 3.1, their approach to detect bundlers requires manually curated fingerprints; They manually analyzed the output of a bundler and gathered unique code snippets which they then used to identify whether a script is a bundled resource. Further, their approach to partially reverse engineer bundles relies on the availability of source maps, i.e., a feature used during development mistakenly being exposed in production. As a result, they were only able to extract modules from around 10% of the bundles they detected in the wild. We instead present an automated approach (see § 3.2 and § 3.3) that can reverse engineer bundles regardless of the accidental availability of source maps. However, in support of our choice of bundler, they found Webpack to be the most popular bundler by an order of magnitude. They observed the presence of 1.27 Webpack bundles on average per site, with 0.11 Browserify bundles per site being the closest alternative choice on the top 100K sites.
@@ 7. Conclusion
Content blocking plays a crucial role in safeguarding privacy, enhancing performance, and preserving user autonomy online. However, the increasing use of bundlers presents a challenge – websites often intermingle tracking code with benign code within a single script – thereby rendering traditional URL-based content blockers ineffective.
We present a framework, Unbundle-Rewrite-Rebundle (URR), that detects bundles and reverse engineers them back to constituent modules. We develop an approach to identify the privacy harming modules that we then replace with benign alternatives. We demonstrate the effectiveness of our system in identifying Webpack bundles and further developed signatures for a fingerprinting library (FingerprintJS), an advertising library (Prebid), and an analytics library (Sentry). Our implementation can identify the bundled versions of these libraries in the wild via a similarity threshold that minimizes false negatives and prevents false positives within our training data. Leveraging our approach, we found the use of these libraries within bundled scripts on 697 sites of the Tranco 10K. Further, we implemented a prototype deployment of URR as a browser extension and observed that the extension only adds 0.5s to the DOM Interaction Time on a page visit.
URR can expand existing content blocking approaches to combat the use of bundlers to hide privacy harming code. It can be further adapted to combat alternative bundling strategies and can be further used to detect the use of other privacy-harming libraries, to bolster web privacy protections.

### v3

@@ 6. Related Work
This research is closely tied to a broad domain of research investigating the advantages, effectiveness, and responses to filter list-based content blocking. Note that existing filter lists (e.g., EasyList, EasyPrivacy) are developed manually and community-maintained. Merzdovnik et al. (Merzdovnik et al., 2017) performed a large-scale study that highlighted the effectiveness of filter list-based browser extensions and also indicated that these tools often lead to a decrease in overall CPU usage, even when considering their own overhead. Gervais et al. (Gervais et al., 2017) quantified the privacy gain from ad-blockers and discovered that these tools can reduce interactions with third-party entities by up to 40% with default settings. However, other studies have pointed out significant inefficiencies in filter lists. Snyder et al. (Snyder et al., 2020) highlighted the abundance of “dead-weight” rules that offer no discernible benefits in popular lists. Similarly, Alrizah et al. (Alrizah et al., 2019) found that popular lists contain a large number of false positives, which can take two or three months to be discovered.
@@
Numerous studies have developed approaches to either help automatically generate filter list rules or develop alternative methods to block privacy-harming content. AdGraph (Iqbal et al., 2020) created a graph-based machine learning approach that used features like URL length and origin to differentiate between benign and privacy-harming resources. Bhagavatula et al. (Bhagavatula et al., 2014) also used URL-based features to train a machine-learning model for resource filtering. Chen et al. (Chen et al., 2021) used filter lists as ground truth and developed behavioral signatures based on the JavaScript event loop, while Sun et al. (Sun et al., 2021) classified JavaScript execution based on Web API calls. Le et al. (Le et al., 2023) developed a reinforcement learning framework that generates filter list rules specific to a site of interest. They showed that their approach was comparable in visual breakage to manually-created filter lists.
@@
Prior work that performs static or dynamic analysis on JavaScript to create “pre-filters” (Fass et al., 2019b) for malicious scripts, detect malicious scripts that camouflage as benign scripts (Fass et al., 2019a), and detect scripts that evade detection by adopting various obfuscation strategies (Fass et al., 2018). In their analysis of the Top 10K sites in the Alexa list, Moog et al. (Moog et al., 2021) found that 90% sites contained a minified or obfuscated script. Fouquet et al. (Fouquet, 2023) presented a static analysis approach to webpack bundles but created signatures based on serializing their AST and hashing the result (compared to the complex, Merkle tree-based approach) we present. Their work does not consider the effect of various bundling and minification options, which our work highlights as highly importance in ensuring the effectiveness of URR as a real-world deployment. Amjad et al. (Amjad et al., 2023) employed stack traces (using dynamic analysis) to identify functions responsible for tracking within scripts. Their work addresses tracking in JavaScript but is not robust against bundles, which include nested functions and commonly introduce code transformations and stack information changes. Importantly, their proposal addresses “tracking behavior”, while URR identified specific libraries. Since our URR is not similarly limited in its ability to identify specific libraries, the benefits provided by the two approaches can be combined. URR can be incorporated as a first step to help narrow the scope of privacy-harming code to a small part of a large, bundled resource. Thereafter,  (Amjad et al., 2023)’s tool can identify specific functions within the library that perform privacy-harming actions, further increasing the granularity of detection and attribution.
More relevant to our work, Rack and Staicu (Rack and Staicu, 2023) presented a method for detecting and partially reverse engineering bundles. Importantly, (Rack and Staicu, 2023) overlaps with the Gather Modules phase (§ 3.1) for which we developed our own approach. To elaborate on the differences between the two approaches, we must divide the Gathering Modules phase into two steps. First, (Rack and Staicu, 2023) determines if a script is a bundle by checking for specific text. Their use of keywords and consideration of a small set of configurations will have limited effect against bundle configurations in the wild. Our approach instead considers code structure and can handle code transformations that would frustrate (Rack and Staicu, 2023). Finally, they did not evaluate their approach’s precision, recall, or accuracy against scripts gathered in the wild (see Table 1). Second, after identifying bundles, (Rack and Staicu, 2023) relies on source maps and unmodified module names to parse bundles and break them down into component modules, restricting their approach to only work on 10% of real-world bundles. We instead present an automated approach (see § 3.2 and § 3.3) that can reverse engineer bundles regardless of the accidental availability of source maps. Our approach does not rely on either of these developer choices and works on a broad range of real-world sites.
@@ 7. Conclusion
Content blocking plays a crucial role in safeguarding privacy, enhancing performance, and preserving user autonomy online. However, the increasing use of bundlers presents a challenge – websites often intermingle tracking code with benign code within a single script, rendering traditional URL-based content blockers ineffective. We present a framework, Unbundle-Rewrite-Rebundle (URR), that detects bundles and reverse engineers them back to constituent modules. We develop an approach to identify the privacy-harming modules that we replace with benign alternatives. We demonstrate the effectiveness of our system in identifying Webpack bundles and further develop signatures for a fingerprinting library (FingerprintJS), an advertising library (Prebid), and an analytics library (Sentry). Our implementation can identify the bundled versions of these libraries in the wild via a similarity threshold that minimizes false negatives and prevents false positives within our training data. Leveraging our approach, we found the use of these libraries within bundled scripts on 697 sites of the Tranco 10K. Further, we implemented a prototype deployment of URR as a browser extension and observed that the extension only adds 0.5s to the DOM Interaction Time on a page visit. URR can expand existing content-blocking approaches to combat the use of bundlers to hide privacy-harming code. It can be further adapted to combat alternative bundling strategies and can be further used to detect the use of other privacy-harming libraries to bolster web privacy protections.
Acknowledgements
We thank Victor Escudero, Marek Cwiek, and Zaheer Safi for performing the manual website analysis. This material is based upon work supported by the National Science Foundation (CNS-2247515).